OpenStudy (anonymous):

Can the expression s times the square root of the quantity 64 times s cubed. + 2stthe square root of s. + 8the square root of the quantity 2 times s. be simplified into an expression with fewer terms? Explain and simplify, if possible.

OpenStudy (campbell_st):

\[s \sqrt{64s^3} + 2s \sqrt{s} + 8\sqrt{2s}\] look at the 1st term \[s \sqrt{64s^3} = s \times \sqrt{64}\times \sqrt{s^2}\times \sqrt{s}\] can you simplify this...?

OpenStudy (anonymous):

well doesnt the square root of 64 turn into 8

OpenStudy (campbell_st):

it does... and there is another part that changes...

OpenStudy (campbell_st):

so you are looking at \[8s \times \sqrt{s^2}\times \sqrt{s}\]

OpenStudy (anonymous):

does the square root of s^7 change

OpenStudy (campbell_st):

what does \[\sqrt{s^2}\] change to..?

OpenStudy (anonymous):

wouldnt it be just s^2

OpenStudy (campbell_st):

no...

OpenStudy (anonymous):

Then i dont know.

OpenStudy (campbell_st):

this is what you told me earlier.. \[\sqrt{64} = \sqrt{8^2} = 8\] use this for \[\sqrt{s^2}\]